OpenStudy (anonymous):

find the derivative of f(x)=(2x^1/4-x^-1/4)^2

OpenStudy (anonymous):

gsrttn do u know how to do this

OpenStudy (anonymous):

Yes I know..

OpenStudy (anonymous):

may you show me how step by step

OpenStudy (anonymous):

\[\large \implies 2[2x^{\frac{1}{4}} - x^{\frac{-1}{4}}][2 \cdot \frac{1}{4} \cdot x^{\frac{-3}{4}} + \frac{1}{4} \cdot x^{\frac{-5}{4}}]\] Can you simplify this by yourself...???

OpenStudy (anonymous):

Do you know the power rule for derivative???

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

are thore brackets supposed to be parenthesis

OpenStudy (richyw):

\[f(x)=(2x^{1/4}-x^{-1/4})^2\]\[\frac{df(x)}{dx}=2\left(2x^{1/4}-x^{-1/4}\right)\frac{d}{dx}\left(2x^{1/4}-x^{-1/4}\right)\]

OpenStudy (richyw):

\[\frac{df(x)}{dx}=2\left(2x^{1/4}-x^{-1/4}\right)\left(\frac{1}{2}x^{-3/4}+\frac{1}{4}x^{-5/4}\right)\]

OpenStudy (richyw):

and some easy but tedious algebra that I don't want to type out gets you to \[\frac{df}{dx}=\frac{4x-1}{2x^{3/2} }\]